OpenStudy (anonymous):

Merchandise was sold by an agent totaling $4,500.00. Her rate of commission was 8%. What was her commission? a.$36.00 b.$360.00 c.$414.00 d.$3,600.00

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

$4500 x .08 = $360

OpenStudy (anonymous):

Which is B.

OpenStudy (gorv):

commision= \[4500*\frac{ 8 }{ 100 }=45*8\]

OpenStudy (anonymous):

Yeah for simple Mental math 4,500/100=45*8=360.